OpenStudy (anonymous):

9a^2/3a+2b-4b^2/3a+2b could someone walk me through this so I can take notes.

OpenStudy (anonymous):

where's the main dash?

OpenStudy (anonymous):

9a^2/3a+2b on bottom then minus 4b^2/3a+2b on bottom

OpenStudy (anonymous):

draw it please

OpenStudy (anonymous):

\[\frac{ 9a^2 }{ 3a+2b? }-\frac{ 4b^2 }{ ?3a+2b }\]

OpenStudy (anonymous):

sorry about the ? marks

OpenStudy (anonymous):

ok ,you see first that both fractions have the same denominator so just subtract the numerators

OpenStudy (anonymous):

so 9-2 =7a^2b^2

OpenStudy (anonymous):

no ,